OpenStudy (anonymous):

find the average rate of change of y=ln (x^2) on the interval (1,2) a. -1.386 b. 0 c. 1.386 d. 2 e. 4

myininaya (myininaya):

\[ARC=\frac{\ln(1^2)-\ln(2^2)}{1-2}\]

myininaya (myininaya):

\[\frac{0-2\ln(2)}{-1}=2\ln(2)\]